OpenStudy (anonymous):

Find the coordinates of the vertex, focus, ends of latus rectum annd the equation of the directrix. a. y^2=8x b. x^2=12y c. y^2=-11x d. x^2=-6y e. (y-5)^2 = 4(x+1) f. (x-7)^2 = -10(y+3)

OpenStudy (anonymous):

how to solve this?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

try to write equation in these form \(y = a(x – h)^2 + k\) where vertex will be \((h,k)\)

OpenStudy (mathmath333):

see if it makes sense

OpenStudy (anonymous):

can we solve letter a. i dont know -.- bzzz..

OpenStudy (mathmath333):

\( y^2=8x\) ,this is a little different from standard quadratic equation , where \(y^2\) is squared instead of \(x\) \(x=\dfrac{y^2}{8}=\dfrac{1}{8}(y-0)^2+0\) here as \((h,k)=(0,0)\) vertex is \((0,0)\)